7. [-/1 Points] DETAILS LARLINALG8 2.3.057. Find A. (2A)-1 = [ 10 -2 1 A = Need Help? Read It Watch It Talk to a Tutor
Answer #1

Given: (2A) = (1 - 4 ► 2A- • det et (12 -1]) = 19-11 = (10)(1)-(-2)(-4) = 10-8=2 adj 15 -1])= (21) adj (12-13) det (C-2-13)